# Deterministic implementation in single-item auctions

### v1

@@ Abstract
Deterministic auctions are attractive in practice due to their transparency, simplicity, and ease of implementation, motivating a sharp understanding of when they match randomized designs. We study deterministic implementation in single-item auctions under two outcome notions: (revenue, welfare) pairs and interim allocations. For (revenue, welfare) pairs, we show a discrete separation: there exists a pair implementable by a deterministic Bayesian incentive-compatible (BIC) auction but not by any deterministic dominant-strategy incentive-compatible (DSIC) auction. For continuous atomless priors, we identify conditions under which deterministic DSIC auctions are implementationally equivalent to randomized BIC auctions in terms of achievable outcomes. For interim allocations, we establish a deterministic analogue of Border’s theorem for two bidders, providing a necessary and sufficient condition for deterministic DSIC implementability, and use it to exhibit an interim allocation implementable by a deterministic BIC auction but not by any deterministic DSIC auction.
@@ 1 Introduction
The central objective of mechanism design is to identify and characterize the set of outcomes that can be implemented by suitable mechanisms. Since the foundational contributions of hurwicz1960optimality; hurwicz1972informationally, a substantial literature has developed across diverse market environments and under multiple solution concepts (clarke1971multipart; dobzinski2022hardness; feldman2023simultaneous; groves1973incentives; myerson1979incentive; myerson1981optimal; myerson1982optimal; vickrey1961counterspeculation; you2021incentive).
This paper advances that agenda in the setting of single-item auctions by studying the implementability of deterministic auctions, where allocation and payment rules are deterministic. Determinism is important in many applications, including antique auctions, spectrum sales, and platform marketplaces, where transparency, compliance, or operational simplicity may rule out randomization (klemperer2004auctions; liu2024optimal). Recent work on single-item auctions shows that, for revenue maximization, optimal auctions can often be chosen deterministic and DSIC in discrete settings, despite allowing randomized BIC mechanisms (giannakopoulos2024discrete). These optimality results, however, leave open a complementary question that is central to implementability: beyond the outcomes induced by optimizing a fixed objective, when does determinism preserve (or fail to preserve) the full set of achievable trade-offs and reduced forms? A precise account of the power and limitations of deterministic auctions is therefore both practically relevant and theoretically significant.
@@ 1.1 Implementation with respect to (revenue, welfare) pairs
We show that deterministic DSIC and deterministic BIC are not equivalent with respect to (revenue, welfare) pairs. There exists a pair that is implementable by a deterministic BIC auction but not by any deterministic DSIC auction. Prior results on revenue inequivalence between DSIC and BIC do not apply here, since they are confined to two-item environments (tang2016optimal; yao2017dominant). To further delineate the boundary of this separation, we identify a restricted class of distributions under which randomized BIC auctions and deterministic DSIC auctions remain equivalent with respect to (revenue, welfare) pairs. In the continuous case, we obtain two levels of equivalence: under regularity we match the Pareto frontier, and under stronger distributional conditions (in the i.i.d. case) we match the entire attainable region.
1.2 Implementation with respect to interim allocation
Our second objective is to characterize which interim allocations can be implemented by deterministic auctions. Here, “interim allocation” refers to the expected allocation probability, taken over the randomness in the other players’ types, in the same sense as defined in myerson1981optimal. border1991implementation gives a necessary and sufficient condition, now called Border’s theorem, for when an interim allocation (the “reduced form” in Border’s terminology) is implementable by a randomized BIC auction. Building on this, manelli2010bayesian show that any interim allocation of a randomized BIC auction is also implementable by a randomized DSIC auction. Further, chen2019equivalence prove that for continuous atomless priors, every interim allocation implementable by a randomized BIC auction is also implementable by a deterministic BIC auction. To the best of our knowledge, no analogous result is known for deterministic DSIC auctions. We show that the deterministic analogue fails: the equivalence does not extend to deterministic DSIC auctions. To show this separation, we introduce new techniques, most notably a deterministic version of Border’s theorem.
@@ 1.3 Our results
We study the power and limitations of deterministic auctions in single-item environments through two outcome notions: (revenue, welfare) pairs and interim allocation.
  • Revenue-welfare outcomes. We characterize which (revenue, welfare) pairs are attainable by deterministic DSIC auctions. In discrete settings, we exhibit a counterexample (via computational enumeration) showing a pair that is implementable by a deterministic BIC auction but by no deterministic DSIC auction, establishing a non-equivalence between the two. In contrast, under continuous regular priors, every Pareto-optimal (revenue, welfare) pair achievable by randomized BIC auctions is implementable by a deterministic DSIC auction. Moreover, in the i.i.d. continuous case, under Condition (1), every (revenue, welfare) pair implementable by randomized BIC auctions is also implementable by a deterministic DSIC auction.
  • Interim allocations. We provide a deterministic analog of Border’s theorem for two bidders, yielding necessary and sufficient conditions for when an interim allocation is implementable by a deterministic DSIC auction. Using this characterization, we show that deterministic BIC and deterministic DSIC are not equivalent. For three bidders with constant interim allocations, we further derive a necessary and sufficient condition for implementability under deterministic DSIC.
1.4 Related works
@@
The study of attainable (revenue, welfare) outcomes is a recurring theme in auction theory (e.g., (bergemann2015limits; ma2022revenue)). hartline2008optimal design an auction that maximizes buyer surplus (welfare minus revenue), which corresponds to an extreme point of the outcome space. pai2009competing analyzes competition among sellers, yielding a revenue-utility trade-off for buyers. kleinberg2013ratio establish lower bounds on the revenue-to-welfare ratio under various feasibility constraints. In a similar spirit, bachrach2014optimising optimize a composite objective $\lambda_{1}$revenue + $\lambda_{2}$welfare + $\lambda_{3}$ click yield” to balance stakeholder goals, and shen2017practical propose a parametric family that optimizes a linear combination of revenue and welfare inside monotone transformations, achieving near-optimal revenue with practical implementability.
@@ 2 Preliminaries
There are $n$ buyers competing for a single item. We use $n]=\{1,2,\cdots,n\}$ to represent the set of $n$ buyers. Each buyer $i$ has a private value $v_{i}\in\mathbb{R}_{\geq 0}$, drawn from a distribution $F_{i}$ with density function $f_{i}$. Valuations are independently drawn across buyers. We use $\boldsymbol{v}=(v_{1},v_{2},\cdots,v_{n})$ to denote the valuation profile of all $n$ buyers and $\boldsymbol{v}_{-i}=(v_{1},\cdots,v_{i-1},v_{i+1},\cdots,v_{n})$ to denote the valuation profile of all buyers except buyer $i$. Each buyer $i$ submits a bid $b_{i}\in\mathbb{R}_{\geq 0}$, and the bid profile is given by $\boldsymbol{b}=(b_{1},b_{2},\cdots,b_{n})$, with $\boldsymbol{b}_{-i}=(b_{1},\cdots,b_{i-1},b_{i+1},\cdots,b_{n})$ representing the bid profile of all buyers except buyer $i$. Throughout, we restrict attention to truthful (incentive-compatible) mechanisms. We use $x_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow[0,1]$, $p_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow\mathbb{R}_{\geq 0}$ to denote the ex-post allocation and the ex-post payment, respectively. An auction is deterministic if $x_{i}(\boldsymbol{v})\in\{0,1\}$ for any $i$ and any $\boldsymbol{v}$. With a slight abuse of notation, we use $x_{i}(v_{i})$, $p_{i}(v_{i})$ to denote interim allocation and interim payment, respectively. The intended meaning will be clear from context. $x_{i}(v_{i})$ and $p_{i}(v_{i})$ are defined as expectations over $\boldsymbol{v}_{-i}$. Specifically, the interim allocation can be expressed as $x_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}x_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$ where $f_{-i}(\boldsymbol{v}_{-i})=\prod_{j\neq i}f_{j}(v_{j})$ and $\mathrm{d}\boldsymbol{v}_{-i}=\mathrm{d}v_{1}\cdots\mathrm{d}v_{i-1}\mathrm{d}v_{i+1}\cdots\mathrm{d}v_{n}$. The interim payment can be expressed as $p_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}p_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$.
We consider the BIC, DSIC, and interim Individually Rational (IR) constraints in mechanisms design. A mechanism is BIC if truth-telling is the optimal strategy for each buyer, given that they have probabilistic beliefs about the valuations of other buyers. A mechanism is DSIC if truth-telling always maximizes a buyer’s utility, regardless of the actions of other buyers. A mechanism is interim IR if each buyer receives a non-negative expected utility from participating. Formally, these constraints are defined as follows:
@@
  • DSIC constraint: $v_{i}\cdot x_{i}(v_{i},\boldsymbol{b}_{-i})-p_{i}(v_{i},\boldsymbol{b}_{-i})\geq v_{i}\cdot x_{i}(b_{i},\boldsymbol{b}_{-i})-p_{i}(b_{i},\boldsymbol{b}_{-i})$, for all $i,v_{i},\boldsymbol{b}_{-i},b_{i}\in\mathbb{R}_{\geq 0}$;
@@
$\textrm{WEL}=\sum_{i=1}^{n}\int_{v_{i}}v_{i}x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},\textrm{REV}=\sum_{i=1}^{n}\int_{v_{i}}\phi_{i}(v_{i})x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},$
where $\phi_{i}(v_{i})=v_{i}-\frac{1-F_{i}(v_{i})}{f_{i}(v_{i})}$, known as the “virtual value” in myerson1981optimal. We also use the quantile space (bulow1989simple). Let $q_{i}=F_{i}(v_{i})\in[0,1]$, and with a slight abuse of notation, let $v_{i}(q_{i})=F_{i}^{-1}(q_{i})$, the inverse function of $F_{i}(v_{i})$. Then, in the quantile space, we have
$\textrm{WEL}=\sum_{i=1}^{n}\int_{q_{i}}v_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},\textrm{REV}=\sum_{i=1}^{n}\int_{q_{i}}\hat{\phi}_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},$
@@
We present a variant of Border’s theorem (border1991implementation; che2013generalized) that, in our setting, characterizes exactly when an interim allocation is implementable by a randomized BIC auction.
@@
$\forall(v_{1},\cdots,v_{n}):~\sum_{i}\int_{t_{i}=v_{i}}^{\infty}x_{i}(t_{i})f_{i}(t_{i})\mathrm{d}{t_{i}}\leq 1-\prod_{i}F_{i}(v_{i}).$
Especially for a symmetric auction in the quantile space, we have
$\forall q_{i}:~\int_{t_{i}=q_{i}}^{1}\hat{x}_{i}(t_{i})\mathrm{d}t_{i}\leq\frac{1-q_{i}^{n}}{n}.$
@@ 3 Implementation with respect to (revenue, welfare) pairs
3.1 Nonequivalence of deterministic BIC auctions and deterministic DSIC auctions with respect to Pareto points in discrete case
In this subsection, we give an example in the i.i.d. discrete case to derive the following theorem.
@@
There exists a Pareto point $P$ that is implemented by a deterministic BIC auction but cannot be implemented by any deterministic DSIC auction in i.i.d discrete case.
Suppose that there are two identical buyers, each with three types: $t_{1}=100,t_{2}=10,t_{3}=1$, occurring with probabilities $p_{1}=0.8,p_{2}=0.15,p_{3}=0.05$ respectively. For each type of buyer 1 and buyer 2, the item can be allocated to no one, to buyer 1, or to buyer 2. Thus, we have to check $3^{9}=19683$ deterministic allocation profiles. Similar to diakonikolas2012efficiency, to consider Pareto-optimal points, we can assume without loss of generality that payments are supported on the prior distribution. Therefore, we can enumerate all deterministic allocation rules on the finite type space and, for each, compute supporting payments and check BIC/DSIC/IR  [^1] Otherwise, there would be infinitely many feasible actions.. Through experiments, we find that $(\textrm{WEL}=96.2275,\textrm{REV}=85)$ is a Pareto-optimal point among all deterministically implementable outcomes (see Figure 1 in Appendix A.1).
@@
The allocation function clearly fails to satisfy the monotonicity requirement for DSIC, as the assignment does not increase monotonically with buyer types. For example, when buyer 1’s type is 1, as buyer 2’s type increases, the allocation transitions from buyer 1 to the seller and then to buyer 2. This transition pattern violates the monotonicity condition required for DSIC implementability. Hence, none of these four auctions can satisfy the dominant-strategy incentive-compatibility constraints, even though they all achieve the same Pareto-optimal (revenue, welfare) pair under BIC.
@@
We prove the theorem in two steps. We first transform the allocation rule into a new one by only changing a single buyer’s interim allocation. We repeat this process $n$ times, and the mechanism with this new allocation rule yields weakly more welfare and revenue under regular distribution condition. In the second step, we show that the last allocation can be implemented in a deterministic DSIC mechanism. Since the original point is already Pareto point of the randomized BIC auction, the transformed auction implements the same point. The proof of Theorem 2 is deferred to Appendix A.2.
@@
The proof sketch is as follows: we first characterize the boundary of randomized BIC auctions, then we prove that each (revenue, welfare) pair on the boundary can be implemented by a deterministic auction (see Lemmas 1 - 2). Finally we prove that each (revenue, welfare) pair strictly between the revenue-maximizing boundary and the revenue-minimizing boundary at the same welfare level can be implemented by a deterministic auction (see Lemma 4). We begin by analyzing the boundary of (revenue, welfare) pairs for randomized BIC auctions. For convenience, we assume that $F$ has full support on $[0,1]$. By (maskin1984optimal), it is without loss of generality to only consider symmetric auctions. Define $P(q)=\int_{q}^{1}\hat{x}(t)\mathrm{d}t$ and we have
$\textrm{WEL}=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}v(q)\mathrm{d}P(q)=n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q,$
$\textrm{REV}=n\int_{0}^{1}\hat{\phi}(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}\hat{\phi}(q)\mathrm{d}P(q)=n\hat{\phi}(0)P(0)+n\int_{0}^{1}\hat{\phi}^{\prime}(q)P(q)\mathrm{d}q.$
@@
$s.t.$ $~~~~n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q=c,$
$~~~~P^{\prime}(q)\leq 0,~P^{\prime\prime}(q)\leq 0,~P(q)\leq\frac{1-q^{n}}{n},~P(1)=0.$

### v2

@@ Abstract
Deterministic auctions are attractive in practice due to their transparency, simplicity, and ease of implementation, motivating a sharper understanding of when they can attain the same outcomes as randomized mechanisms. We study deterministic implementation in single-item auctions under two notions of outcomes: (revenue, welfare) pairs and interim allocations. For (revenue, welfare) pairs, we show a separation in discrete settings: there exists a pair implementable by a deterministic Bayesian incentive-compatible (BIC) auction but not by any deterministic dominant-strategy incentive-compatible (DSIC) auction. For continuous atomless priors, we identify conditions under which deterministic DSIC auctions are equivalent to randomized BIC auctions in terms of achievable outcomes. For interim allocations, under a strict monotonicity condition, we establish a deterministic analogue of Border’s theorem for two bidders, providing a necessary and sufficient condition for deterministic DSIC implementability. Using this characterization, we exhibit an interim allocation implementable by a randomized BIC auction but not by any deterministic DSIC auction.
@@ 1 Introduction
The central objective of mechanism design is to identify and characterize the set of outcomes that can be implemented by suitable mechanisms. Since the foundational contributions of hurwicz1960optimality; hurwicz1972informationally, a substantial literature has developed across diverse market environments and under multiple solution concepts clarke1971multipart; dobzinski2022hardness; feldman2023simultaneous; groves1973incentives; myerson1979incentive; myerson1981optimal; myerson1982optimal; vickrey1961counterspeculation; you2021incentive.
This paper advances that agenda in the setting of single-item auctions by studying the implementability of deterministic auctions, where allocation and payment rules are deterministic. Determinism is important in many applications, including antique auctions, spectrum sales, and platform marketplaces, where transparency, compliance, or operational simplicity may rule out randomization klemperer2004auctions; liu2024optimal. Recent work on single-item auctions shows that, for revenue maximization, optimal auctions can often be chosen as deterministic DSIC mechanisms in discrete settings, despite allowing randomized BIC mechanisms giannakopoulos2024discrete. These optimality results, however, leave open a complementary question that is central to implementability: beyond the outcomes induced by optimizing a fixed objective, when does determinism preserve (or fail to preserve) the full set of achievable trade-offs and reduced forms? A precise account of the power and limitations of deterministic auctions is therefore both practically relevant and theoretically significant.
@@ 1.1 Implementation with Respect to (Revenue, Welfare) Pairs
We show that deterministic DSIC and deterministic BIC are not equivalent with respect to (revenue, welfare) pairs. There exists a pair that is implementable by a deterministic BIC auction but not by any deterministic DSIC auction. Prior results on revenue inequivalence between DSIC and BIC do not apply here, since they are confined to two-item environments tang2016optimal; yao2017dominant. To further delineate the boundary of this separation, we identify a restricted class of distributions under which randomized BIC auctions and deterministic DSIC auctions remain equivalent with respect to (revenue, welfare) pairs. In the continuous case, we obtain two levels of equivalence: under regularity we match the Pareto frontier, and under stronger distributional conditions (in the i.i.d. case) we match the entire attainable region.
1.2 Implementation with Respect to Interim Allocations
Our second objective is to characterize which interim allocations can be implemented by deterministic auctions. Here, “interim allocation” refers to the expected allocation probability, taken over the randomness in the other players’ types, in the same sense as defined in myerson1981optimal. border1991implementation gives a necessary and sufficient condition, now called Border’s theorem, for when an interim allocation (the “reduced form” in Border’s terminology) is implementable by a randomized BIC auction. Building on this, manelli2010bayesian show that any interim allocation of a randomized BIC auction is also implementable by a randomized DSIC auction. Further, chen2019equivalence prove that for continuous atomless priors, every interim allocation implementable by a randomized BIC auction is also implementable by a deterministic BIC auction. To the best of our knowledge, no analogous result is known for deterministic DSIC auctions. We show that an analogous equivalence fails for deterministic DSIC auctions. To show this separation, we introduce new techniques, most notably a deterministic version of Border’s theorem.
@@ 1.3 Our Results
We study the power and limitations of deterministic auctions in single-item environments through two outcome notions: (revenue, welfare) pairs and interim allocations.
  • Revenue-welfare outcomes. We characterize which (revenue, welfare) pairs are attainable by deterministic DSIC auctions. In discrete settings, we exhibit a counterexample (via computational enumeration) showing a pair that is implementable by a deterministic BIC auction but not by any deterministic DSIC auction, establishing a non-equivalence between the two. In contrast, under continuous regular priors, every Pareto-optimal (revenue, welfare) pair achievable by randomized BIC auctions is implementable by a deterministic DSIC auction. Moreover, in the i.i.d. continuous case, under Condition (1), every (revenue, welfare) pair implementable by randomized BIC auctions is also implementable by a deterministic DSIC auction.
  • Interim allocations. We provide a deterministic analogue of Border’s theorem for two bidders under a strict monotonicity condition, yielding a necessary and sufficient condition under which an interim allocation is implementable by a deterministic DSIC auction. Using this characterization, we show that randomized BIC and deterministic DSIC auctions are not equivalent. For three bidders with constant interim allocations, we further derive a necessary and sufficient condition for implementability under deterministic DSIC.
1.4 Related Work
@@
The study of attainable (revenue, welfare) outcomes is a recurring theme in auction theory (e.g., bergemann2015limits; ma2022revenue). hartline2008optimal design an auction that maximizes buyer surplus (welfare minus revenue), which corresponds to an extreme point of the outcome space. pai2009competing analyzes competition among sellers, yielding a revenue-utility trade-off for buyers. kleinberg2013ratio establish lower bounds on the revenue-to-welfare ratio under various feasibility constraints. In a similar spirit, bachrach2014optimising optimize a composite objective $\lambda_{1}\cdot\mathrm{revenue}$ + $\lambda_{2}\cdot\mathrm{welfare}$ + $\lambda_{3}\cdot\mathrm{click~yield}$” to balance stakeholder goals, and shen2017practical propose a parametric family that optimizes a linear combination of revenue and welfare inside monotone transformations, achieving near-optimal revenue with practical implementability.
@@ 2 Preliminaries
There are $n$ buyers competing for a single indivisible item. Let $n]=\{1,2,\cdots,n\}$ represent the set of buyers. Buyer $i$ has a private valuation $v_{i}\in\mathbb{R}_{\geq 0}$, drawn independently from a (not necessarily identically distributed) prior distribution $F_{i}$ with density function $f_{i}$. We use $\boldsymbol{v}=(v_{1},v_{2},\cdots,v_{n})$ to denote the valuation profile of all $n$ buyers and $\boldsymbol{v}_{-i}=(v_{1},\cdots,v_{i-1},v_{i+1},\cdots,v_{n})$ to denote the valuation profile of all buyers except buyer $i$. Each buyer $i$ submits a bid $b_{i}\in\mathbb{R}_{\geq 0}$, and the bid profile is given by $\boldsymbol{b}=(b_{1},b_{2},\cdots,b_{n})$, with $\boldsymbol{b}_{-i}=(b_{1},\cdots,b_{i-1},b_{i+1},\cdots,b_{n})$ representing the bid profile of all buyers except buyer $i$. Throughout, we restrict attention to truthful (incentive-compatible) mechanisms. We use $x_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow[0,1]$, $p_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow\mathbb{R}_{\geq 0}$ to denote the ex-post allocation and the ex-post payment, respectively. An auction is deterministic if $x_{i}(\boldsymbol{v})\in\{0,1\}$ for any $i$ and any $\boldsymbol{v}$. With a slight abuse of notation, we use $x_{i}(v_{i})$, $p_{i}(v_{i})$ to denote interim allocation and interim payment, respectively. The intended meaning will be clear from context. $x_{i}(v_{i})$ and $p_{i}(v_{i})$ are defined as expectations over $\boldsymbol{v}_{-i}$. Specifically, the interim allocation can be expressed as $x_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}x_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$ where $f_{-i}(\boldsymbol{v}_{-i})=\prod_{j\neq i}f_{j}(v_{j})$ and $\mathrm{d}\boldsymbol{v}_{-i}=\mathrm{d}v_{1}\cdots\mathrm{d}v_{i-1}\mathrm{d}v_{i+1}\cdots\mathrm{d}v_{n}$. The interim payment can be expressed as $p_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}p_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$.
We consider the BIC, DSIC, and interim Individually Rational (IR) constraints in mechanism design. A mechanism is BIC if truth-telling is the optimal strategy for each buyer, given that they have probabilistic beliefs about the valuations of other buyers. A mechanism is DSIC if truth-telling always maximizes a buyer’s utility, regardless of the actions of other buyers. A mechanism is interim IR if each buyer receives a non-negative expected utility from participating. Formally, these constraints are defined as follows:
@@
  • DSIC constraint: $v_{i}\cdot x_{i}(v_{i},\boldsymbol{b}_{-i})-p_{i}(v_{i},\boldsymbol{b}_{-i})\geq v_{i}\cdot x_{i}(b_{i},\boldsymbol{b}_{-i})-p_{i}(b_{i},\boldsymbol{b}_{-i})$, for all $i$, $v_{i}$, $b_{i}\in\mathbb{R}_{\geq 0}$, and $\boldsymbol{b}_{-i}\in\mathbb{R}_{\geq 0}^{n-1}$;
@@
WEL $=\sum_{i=1}^{n}\int_{v_{i}}v_{i}x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},$
REV $=\sum_{i=1}^{n}\int_{v_{i}}\phi_{i}(v_{i})x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},$
where $\phi_{i}(v_{i})=v_{i}-\frac{1-F_{i}(v_{i})}{f_{i}(v_{i})}$, known as the “virtual value” in myerson1981optimal. We also use the quantile space bulow1989simple. Let $q_{i}=F_{i}(v_{i})\in[0,1]$, and with a slight abuse of notation, let $v_{i}(q_{i})=F_{i}^{-1}(q_{i})$, the inverse function of $F_{i}(v_{i})$. Then, in the quantile space, we have
WEL $=\sum_{i=1}^{n}\int_{q_{i}}v_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},$
REV $=\sum_{i=1}^{n}\int_{q_{i}}\hat{\phi}_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},$
@@
We present a variant of Border’s theorem border1991implementation; che2013generalized that, in our setting, characterizes exactly when an interim allocation is implemented by a randomized BIC auction.
@@
$\forall(v_{1},\cdots,v_{n}):\sum_{i}\int_{t_{i}=v_{i}}^{\infty}x_{i}(t_{i})f_{i}(t_{i})\mathrm{d}{t_{i}}\leq 1-\prod_{i}F_{i}(v_{i}).$
In the i.i.d. symmetric case, in quantile space, it simplifies to
$\forall q\in[0,1]:\int_{q}^{1}\hat{x}(t)\mathrm{d}t\leq\frac{1-q^{n}}{n},$
where $\hat{x}$ denotes the common interim allocation function.
The inequality compares two probabilities that must be consistent in any feasible single-item allocation. Fix thresholds $(v_{1},\cdots,v_{n})$ and consider the event that every bidder $i$’s type exceeds $v_{i}$. The left-hand side, $\sum_{i}\int_{t_{i}=v_{i}}^{\infty}x_{i}(t_{i})f_{i}(t_{i})\mathrm{d}t_{i}$, is the ex-ante probability that the mechanism allocates the item to some bidder from within the set of “above-threshold” types (counting bidder-by-bidder). The right-hand side, $1-\prod_{i}F_{i}(v_{i})$, is the probability that at least one bidder is above her threshold. Since only one item can be allocated, the probability of allocating the item to an above-threshold bidder can never exceed the probability that such a bidder exists. Border’s theorem states that this family of inequalities (for all thresholds) is not only necessary but also sufficient for interim implementability. In our paper, this characterization lets us work directly with interim allocation rules in value/quantile space: we propose candidate interim allocations $\hat{x}$, verify feasibility by checking the corresponding Border constraints (e.g., the symmetric quantile form $\int_{q}^{1}\hat{x}(t)\mathrm{d}t\leq\frac{1-q^{n}}{n}$), and then optimize linear objectives such as welfare and (virtual) revenue over this feasible set. Finally, tightness of a Border constraint has a clean meaning: equality at $(v_{1},\cdots,v_{n})$ means there is no slack in the event “someone is above threshold”, i.e., whenever at least one bidder exceeds her threshold, the mechanism allocates the item with probability 1 to an above-threshold bidder. Strict inequality indicates that with positive probability the mechanism leaves allocation probability “unused” relative to that event (for example, it sometimes does not allocate to any above-threshold bidder even though one exists). This interpretation is key for understanding the randomized benchmark region and for contrasting it with the additional structure imposed by deterministic DSIC implementability later in the paper.
@@ 3 Implementation with Respect to (Revenue, Welfare) Pairs
3.1 Nonequivalence of Deterministic BIC and Deterministic DSIC Auctions with Respect to Pareto Points in the Discrete Case
In Section 2, values are drawn from continuous distributions. Here we consider an i.i.d. discrete prior: each $v_{i}$ takes values in $V=\{v^{1},\cdots,v^{m}\}$ with probability vector $\boldsymbol{\pi}=(\pi_{1},\cdots,\pi_{m})$, where $\pi_{k}=\mathrm{Pr}[v_{i}=v^{k}]$. A deterministic mechanism $(x_{i},p_{i})$ specifies an ex post allocation $x_{i}(\mathbf{v})\in\{0,1\}$ and payment $p_{i}(\mathbf{v})$ for every profile $\mathbf{v}\in V^{n}$. We evaluate it by expected welfare and expected revenue under the prior $\boldsymbol{\pi}$:
$\mathrm{WEL}$ $=\mathbb{E}_{\mathbf{v}\sim\boldsymbol{\pi}^{n}}\Big[\sum_{i\in[n]}v_{i}x_{i}(\mathbf{v})\Big],$
$\mathrm{REV}$ $=\mathbb{E}_{\mathbf{v}\sim\boldsymbol{\pi}^{n}}\Big[\sum_{i\in[n]}p_{i}(\mathbf{v})\Big].$
Equivalently, $\mathbb{E}_{\mathbf{v}\sim\boldsymbol{\pi}^{n}}[\cdot]$ denotes a finite sum over all $\mathbf{v}=(v_{1},\cdots,v_{n})\in V^{n}$, where each coordinate $v_{i}=v^{k_{i}}$ is weighted by $\pi_{k_{i}}$. Hence the profile $\mathbf{v}$ has weight $\prod_{i\in[n]}\pi_{k_{i}}$.
We give an example in the i.i.d. discrete case to derive the following theorem.
@@
There exists a Pareto point $P$ that is implemented by a deterministic BIC auction but cannot be implemented by any deterministic DSIC auction in the i.i.d. discrete case.
Suppose that there are two identical buyers, each with three types: $t_{1}=100,t_{2}=10,t_{3}=1$, occurring with probabilities $p_{1}=0.8,p_{2}=0.15,p_{3}=0.05$ respectively. For each type of buyer 1 and buyer 2, the item can be allocated to no one, to buyer 1, or to buyer 2. Thus, we have to check $3^{9}=19683$ deterministic allocation profiles. Similar to diakonikolas2012efficiency, to consider Pareto-optimal points, we can assume without loss of generality that payments are specified on the finite type space. Therefore, we can enumerate all deterministic allocation rules on the finite type space and, for each, compute supporting payments and check BIC/DSIC/IR  [^1] Otherwise, there would be infinitely many feasible actions.. Through experiments, we find that $(\textrm{WEL}=96.2275,\textrm{REV}=85)$ is a Pareto-optimal point among all outcomes implementable by deterministic BIC mechanisms (see Figure 1). Only partial figure is provided here since the complete figure is too large.
@@
Although the interim allocation rules are monotone (hence the mechanism is BIC), the ex-post allocation rule is not monotone in a bidder’s own report for fixed opponent reports, so it cannot be DSIC. For instance, fixing buyer 2’s report at $v_{2}=1$, buyer 1 wins when $v_{1}=1$, loses when $v_{1}=10$, and wins again when $v_{1}=100$, which violates the monotonicity requirement for deterministic DSIC implementability. Hence, none of these four auctions can satisfy the DSIC constraints, even though they all achieve the same Pareto-optimal (revenue, welfare) pair under BIC.
@@
We prove the theorem in two steps. In the first step, starting from a randomized BIC mechanism, we transform its allocation rule by modifying only one buyer’s interim allocation at a time, while keeping the other buyers’ interim allocations fixed. We repeat this procedure for all $n$ buyers. Under the regularity condition, each transformation weakly improves both expected welfare and expected revenue, so the final mechanism weakly dominates the original one in the $(\mathrm{REV},\mathrm{WEL})$ space. In the second step, we show that the last (transformed) allocation rule admits a deterministic DSIC implementation. Since the original point is already Pareto-optimal among outcomes achievable by randomized BIC mechanisms, the weak improvement must be tight. Therefore, the deterministic DSIC mechanism implements the same Pareto point. The full proof of Theorem 2 is deferred to Appendix A.1.
@@
The proof sketch is as follows: we first characterize the boundary of randomized BIC auctions, then we prove that each (revenue, welfare) pair on the boundary can be implemented by a deterministic auction (see Lemmas 1 - 2). Finally we prove that each (revenue, welfare) pair strictly between the revenue-maximizing boundary and the revenue-minimizing boundary at the same welfare level can be implemented by a deterministic auction (see Lemma 4). The full proof of Theorem 3 is deferred to Appendix A.6.
We begin by analyzing the boundary of (revenue, welfare) pairs for randomized BIC auctions. For convenience, we assume that $F$ has full support on $[0,1]$. By maskin1984optimal, it is without loss of generality to only consider symmetric auctions. Define $P(q)=\int_{q}^{1}\hat{x}(t)\mathrm{d}t$ and we have
WEL $=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}v(q)\mathrm{d}P(q)$
$=n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q,$
REV $=n\int_{0}^{1}\hat{\phi}(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}\hat{\phi}(q)\mathrm{d}P(q)$
$=n\hat{\phi}(0)P(0)+n\int_{0}^{1}\hat{\phi}^{\prime}(q)P(q)\mathrm{d}q.$
@@
s.t. $n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q=c,~P(q)\leq\frac{1-q^{n}}{n},$
$P^{\prime}(q)\leq 0,~P^{\prime\prime}(q)\leq 0,~P(1)=0.$